Phenyl acetylene reacts with dl. H2SO4 in presence of HgSO4 gives?

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
Dear Yogiraj,
Phenyl acetylene + H2SO4 + HgSO4 → (C6H5)- (C-OH)=CH → C6H5-(C=O) – CH3
Final Product  will be C6H5-(C=O) – CH3